# ResuméAtlas: Revisiting Resume Classification with Large-Scale Datasets and Large Language Models

### v1

@@ Abstract
The increasing reliance on online recruitment platforms coupled with the adoption of AI technologies has highlighted the critical need for efficient resume classification methods. However, challenges such as small datasets, lack of standardized resume templates, and privacy concerns hinder the accuracy and effectiveness of existing classification models. In this work, we address these challenges by presenting a comprehensive approach to resume classification. We curated a large-scale dataset of 13,389 resumes from diverse sources and employed Large Language Models (LLMs) such as BERT and Gemma1.1 2B for classification. Our results demonstrate significant improvements over traditional machine learning approaches, with our best model achieving a top-1 accuracy of 92% and a top-5 accuracy of 97.5%. These findings underscore the importance of dataset quality and advanced model architectures in enhancing the accuracy and robustness of resume classification systems, thus advancing the field of online recruitment practices.
@@ 5 Results & Discussion
6 Conclusion
In this study, we addressed the significant challenges in resume classification within the online recruitment domain by leveraging a large-scale dataset and advanced Language Model Models (LLMs). Through meticulous dataset collection efforts, we assembled a comprehensive dataset comprising 13,389 resumes across 43 distinct categories. This extensive dataset enabled us to overcome limitations associated with small datasets and variability in resume formats, thereby enhancing the robustness and reliability of our classification models. Additionally, by employing state-of-the-art transformer models such as BERT and Gemma1.1 2B, we achieved remarkable improvements in classification accuracy, outperforming traditional machine learning approaches.
Our results demonstrate the efficacy of utilizing large datasets and advanced LLMs in improving the accuracy of resume classification tasks. The top-performing models achieved top-1 accuracy of 92% and a top-5 accuracy of 97.5%, underscoring the effectiveness of our approach in addressing the complexities of resume classification.
7 Future work
While our study represents a significant advancement in the field of resume classification, there are several avenues for future research and improvement. One promising direction is the collection of even larger datasets from a diverse range of sources, including social media platforms, job boards, and professional networking sites. Additionally, expanding the scope of labels and job titles covered in the dataset can further enhance the generalization capabilities of classification models. Furthermore, exploring innovative techniques for handling resume variability, such as adapting to different resume formats and incorporating multi-modal information, holds promise for improving classification accuracy and robustness. Overall, continued efforts in dataset collection, model development, and methodological innovation are essential for advancing the state-of-the-art in resume classification and enhancing the efficiency of online recruitment practices.

### v2

@@ Abstract
The increasing reliance on online recruitment platforms coupled with the adoption of AI technologies has highlighted the critical need for efficient resume classification methods. However, challenges such as small datasets, lack of standardized resume templates, and privacy concerns hinder the accuracy and effectiveness of existing classification models. In this work, we address these challenges by presenting a comprehensive approach to resume classification. We curated a large-scale dataset of 13,389 resumes from diverse sources and employed Large Language Models (LLMs) such as BERT and Gemma1.1 2B for classification. Our results demonstrate significant improvements over traditional machine learning approaches, with our best model achieving a top-1 accuracy of 92% and a top-5 accuracy of 97.5%. These findings underscore the importance of dataset quality and advanced model architectures in enhancing the accuracy and robustness of resume classification systems, thus advancing the field of online recruitment practices. Our models, code, and dataset are available as open-source resources.  [^2] Code: https://github.com/noran-mohamed/Resume-Classification-Dataset  [^3] Models: https://huggingface.co/collections/ahmedheakl/resumeatlas-668047e86bc332049afd0b39  [^4] Dataset: https://huggingface.co/datasets/ahmedheakl/resume-atlas
@@ 1 Introduction
The rest of this paper is presented as follows: Section 2 reviews related works in the field of resume classification. Section 3 describes our dataset collection process, challenges, and preprocessing steps. Section 4 outlines the methodologies employed in our study, including the use of LLMs and traditional machine learning approaches. Section 5 presents and discusses our experimental results. Finally, Section 6 concludes the paper and suggests directions for future research.
@@ 5 Results & Discussion
6 Conclusion and Future Work
In this study, we addressed significant challenges in resume classification within the online recruitment domain by leveraging a large-scale dataset and advanced Language Model Models (LLMs). Through meticulous dataset collection efforts, we assembled a comprehensive dataset of 13,389 resumes across 43 distinct categories, overcoming limitations associated with small datasets and resume format variability. By employing state-of-the-art transformer models such as BERT and Gemma1.1 2B, we achieved remarkable improvements in classification accuracy, with top-performing models reaching 92% top-1 accuracy and 97.5% top-5 accuracy, outperforming traditional machine learning approaches.
Future research directions include expanding the dataset with diverse sources, broadening the scope of job titles, and exploring techniques for handling resume variability. Investigating domain-specific pre-training for LLMs and developing methods to interpret model decisions could further enhance the applicability and transparency of resume classification in real-world recruitment scenarios.